# Less is More: Mitigating Multimodal Hallucination from an EOS Decision Perspective

### v1

@@ Abstract
Large Multimodal Models (LMMs) often suffer from multimodal hallucinations, wherein they may create content that is not present in the visual inputs. In this paper, we explore a new angle of this issue: overly detailed training data hinders the model’s ability to timely terminate generation, leading to continued outputs beyond visual perception limits. By investigating how the model decides to terminate generation with EOS, the special end-of-sentence token, we find that the model assesses the completeness of the entire sequence by comparing the generated text with the image. This observation suggests that the model possesses an inherent potential of making proper EOS decisions based on its visual perception to avoid overly lengthy outputs. To take advantage of such potential, we explore two methods to mitigate multimodal hallucinations: a training objective that enables the model to reduce hallucinations by learning from regular instruction data, and a data filtering strategy to prevent harmful training data from exacerbating model hallucinations. Both methods significantly improve the hallucination performance of LMMs, without requiring any additional data or knowledge.111https://github.com/yuezih/less-is-more
@@
Zihao Yue Renmin University of China yzihao@ruc.edu.cn Liang Zhang Renmin University of China zhangliang00@ruc.edu.cn Qin Jin Renmin University of China qjin@ruc.edu.cn
@@ 1 Introduction
Ever since Large Multimodal Models (LMMs) Yin et al. (2023a) were achieved through bridging vision encoders with Large Language Models (LLMs) Brown et al. (2020); Zhao et al. (2023a), they have been plagued by the problem of multimodal hallucinations, i.e., their text outputs may include unfaithful content to the visual inputs, such as non-existent objects Rohrbach et al. (2018), which greatly harms the reliability of their applications. Extensive research has shed light on the origins of multimodal hallucinations, including the inability of vision encoders to represent fine-grained visual details Jiang et al. (2023b); Tong et al. (2024), model reliance on inherent parametric knowledge such as language priors and statistical biases Leng et al. (2023); Zhou et al. (2023), and pervasive hallucinations in the training data itself Yu et al. (2023a); Liu et al. (2023a). In response to these insights, a variety of strategies have been proposed to mitigate hallucinations in LMMs Yin et al. (2023b); Zhai et al. (2023); Zhao et al. (2023b); Yu et al. (2023b).
Although significant progress has been made, in this paper, we highlight a crucial but often overlooked source of hallucinations: the excessively detailed training data. For example, in the detailed image captioning task, the caption data for an image typically integrates rich visual semantics from multiple human annotations or vision expert models, and is rewritten into lengthy paragraphs by LLMs Liu et al. (2023c), as shown in Fig. 1. These training data, while high-quality and meeting our expectations for detail, may exceed the visual perception capability of LMMs, especially for subtle image features such as small or easily confusable objects. When trained with such data, in an attempt to fit the detail level and length distribution of ground truth captions, the model may risk expressing details that it cannot discern from the image, and therefore exhibit hallucinations.
@@
To unlock such potential of models, we explore two approaches to enhance the model for better EOS decisions. (1) A learning objective for model training, termed Selective EOS Supervision. Simply modified from the Maximum Likelihood Estimation (MLE), it enables the model to mitigate hallucinations through learning from regular instruction data. It is applicable both for further training to reduce hallucinations in existing models, and for initial instruction tuning Ouyang et al. (2022) to alleviate the onset of hallucinations. Specifically, by briefly further training on the original instruction data, the sentence-level and instance-level hallucinations of LLaVA-1.5 Liu et al. (2023b) are reduced by 26% and 27%, respectively. (2) A data filtering strategy based on Scoring EOS Supervision, to eliminate harmful training data that can impair the model’s ability to end sequences. We design two metrics to assess the positive and negative impact of data on the model’s EOS tendency, and combine them to rank and filter the training data. Experimental results show that removing a small portion of the data can significantly reduce the hallucinations of models trained on it. These findings further validate our hypothesis and provide simple yet effective solutions for mitigating multimodal hallucinations in LMMs.
@@ 2 EOS Decision
In this section, we delve into how LMMs reach EOS decisions. Specifically, in Section 2.1, we analyze the contextual information that the model relies on to predict ; in Section 2.2, we explore how the model adjusts its tendency to terminate generation with the multimodal input. Corresponding findings are further discussed in Section 2.3.
@@ 2.1 Information Basis of EOS Decision
where denotes the self-attention matrix, and reflects the significance of the information flow from the -th token to the -th token. We compare the information flow patterns of EOS predictions and non-EOS predictions to elucidate the information basis of EOS decisions.
@@ 3.1 Selective EOS Supervision for Training
The instruction tuning of LMMs typically utilizes Maximum Likelihood Estimation (MLE) as the training objective. Given the visual content and previous tokens , the model predicts a probability distribution over the vocabulary to determine the next word, where represents the probability of the -th word in . The model parameter is optimized to maximize the likelihood of the label word indexed , with the loss function defined as:
@@
Row | Model | Method | Length | CHAIR | CHAIR | Recall | F-Score | F-Score |
@@
5 | Truncate (length) | 79.9 | 41.8 | 13.2 | 72.5 | 87.8 | 69.8 | |
6 | Truncate (proportion) | 79.8 | 43.4 | 14.0 | 74.3 | 88.2 | 72.1 | |
7 | Truncate (sentence) | 80.0 | 42.8 | 14.1 | 74.2 | 87.9 | 71.6 | |
@@ 3.1.1 Experimental Settings
Models and datasets. Our training objective can be applied to any LMMs with an EOS token and optimized by MLE. As a representative, we select two widely used open-source LMMs, LLaVA Liu et al. (2023c) and MiniGPT Zhu et al. (2023) series. Among them, LLaVA, LLaVA-1.5 Liu et al. (2023b), and MiniGPT-v2 Chen et al. (2023a) are trained with data recipes that include the LLaVA-Instruction dataset. Thus, we validate our method with these models by fine-tuning them on LLaVA-Instruction. Additionally, our experiment results show that a smaller subset of LLaVA-Instruction-150K which contains detailed captions, Detail23K, has a similar effect but brings significant computational efficiency. Thus, most of our experiments are conducted with Detail23K. If not specified, models undergo one epoch of training with LoRA Hu et al. (2022). Other training details remain consistent with the models’ official documentation.
@@
Evaluation. Following previous works Huang et al. (2023); Yu et al. (2023a), we evaluate model hallucination with Caption Hallucination Assessment with Image Relevance (CHAIR) Rohrbach et al. (2018), a framework that quantifies object hallucination in image captions by comparing generated objects to the ground truth objects. The sentence-level score, CHAIR, represents the proportion of captions that contain hallucinations, and the instance-level score, CHAIR, denotes the frequency of hallucinated objects relative to all mentioned objects by the model. In addition, we measure an object recall to evaluate the semantic richness of generated captions. Our CHAIR tests involve 500 images randomly chosen from the MSCOCO validation set Lin et al. (2014). We also adopt another metric FaithScore Jing et al. (2023) to evaluate caption hallucination. It verifies the consistency between atomic facts in the caption and the input image with LLMs and visual expert models, for which we employ ChatGPT OpenAI (2023) and OFA Wang et al. (2022). It also provides a sentence-level score, FaithScore.
@@
Baselines. Since our method facilitates models to timely terminate generation, we design three baselines for comparison by simply truncating the generation results. (1) Truncate by length: We set a global length maximum , and truncate captions that exceed this length; (2) Truncate by proportion: We define a proportion , and keep only the initial of words in each caption. (3) Truncate by sentence: We remove the last sentences from each caption. The parameters , , and are adjusted to obtain similar average caption lengths for fair comparison. Additionally, we include two recently proposed plug-in methods: (1) Visual Contrastive Decoding (VCD) Leng et al. (2023), which contrasts the output distributions derived from the original and noisy visual inputs, to reduce the influence of the model’s parametric knowledge. (2) Over-Trust Penalty and Retrospection-Allocation (OPERA) Huang et al. (2023), a decoding strategy that penalizes the model’s over-reliance on certain tokens and allows roll-back when needed. We test VCD at different noise steps of 200, 500, 700, and 999, and report the optimal results. For OPERA, our implementation follows the suggestions in their released code, including two hyperparameter configurations for standard and fast inference.
@@ 3.1.2 Results
Versus the original model. As shown in Table 1, after a single training epoch on the detailed caption subset, Detail23K, using our learning objective, all models tend to produce shorter paragraphs and notably reduce hallucinations at both the sentence and instance levels ( vs , vs , etc.). This improvement is even more significant when using the full 150K instruction data ( vs ), resulting in a 26.4% and 26.6% decrease in CHAIR and CHAIR of LLaVA-1.5 (7b), respectively. While our method leads to some decrease in recall (e.g., 1.8% of LLaVA-1.5 (7b) and 2.6% of LLaVA-1.5 (13b)), we view this as a beneficial compromise since the models become more conservative and less likely to “guess” uncertain visual content.
Versus baselines. As shown in , , and of Table 1, our designed three naive Truncate methods effectively reduce hallucinations in the model outputs, indicating that while introducing incoherence, simply shortening the sequence length can serve as strong baselines for mitigating hallucinations. Nonetheless, our method performs even better, particularly excelling these baselines on CHAIR by a large margin and incurring less recall loss. It also outperforms existing methods, i.e., VCD () and OPERA (, ). Our proposed method does not require additional data construction or other expert models. Furthermore, unlike decoding methods, it does not slow down inference and remains technically compatible with various decoding strategies. Therefore, it presents a viable, practical supplement or alternative to current methods.
@@
Loss | Length | CHAIR | CHAIR | Recall |
@@ 4 Related Works
Hallucination Origins. Investigations on the causes of multimodal hallucination in LMMs identify three main factors: (1) Limited visual representations. For example, the visual encoders commonly employed in LMMs depict abstract features while struggling to capture fine-grained visual details Jiang et al. (2023b); Tong et al. (2024); Wang et al. (2023a). Jiang et al. (2023a) observe a modality gap persists between visual and textual features, despite vision-language alignment. (2) Models’ over-reliance on parametric knowledge, such as statistical biases and language priors, rather than on visual evidence Zhou et al. (2023); Leng et al. (2023); Liu et al. (2023a); Zhai et al. (2023); Guan et al. (2023). (3) Inferior data for instruction tuning. This includes insufficient visual supervision Chen et al. (2023b), a lack of positive/negative human feedback Yu et al. (2023b), and the presence of hallucinations within the training data Yu et al. (2023a); Liu et al. (2023a). Our paper identifies a new source of hallucinations: overly detailed training data hinders the model’s inherent EOS decision ability, further enriching existing explanations.
@@ 5 Conclusion
This work focuses on reducing hallucinations in large multimodal models to enhance their reliability and trustworthiness. We have carefully considered the ethical implications of our work and anticipate no significant ethical concerns. This work was carried out using publicly available and commonly-used data and models, and our findings may inherit the biases and limitations carried in these resources.
@@ B.2 Context Manipulation
B.3 EOS Prediction Tendency
In Fig. 8, we illustrate the EOS prediction tendency (average probability) of the LLaVA-1.5 (7b) model during further training on Detail23K. With Selective EOS Supervision proposed in Section 3.1, the model’s tendency to predict EOS rises and stabilizes, while the model optimized by MLE shows no change in this behavior. This suggests that the proposed training objective effectively helps the model regain its capability to timely conclude sequences.
B.4 Data Score Distributions
@@
LLaVA-1.5 (7b) | Baseline | 1516.1 | 348.2 | 85.9 | 86.9 | 94.0 | 79.1 |
MLE | 1470.9 | 372.5 | 86.1 | 87.0 | 93.6 | 79.7 | |
Ours | 1490.4 | 367.9 | 86.0 | 86.8 | 93.5 | 79.5 | |
@@
In Section 3.2, we discuss two metrics, and , which are summed over positions labeled as EOS and non-EOS, respectively. A natural concern is that the number of non-EOS positions far exceeds that of EOS positions, raising the question of whether the combined might be dominated by . To clarify this, we examine the score distributions within the LLaVA-Instruction-150K dataset. As illustrated in Fig. 9, the value distributions of and are comparable in magnitude, and the distribution is approximately normal with zero mean. This indicates a balance between and with neither metric dominating, and the top scores necessitate both high and low . Thus, by maintaining a straightforward formulation of without introducing a balancing hyperparameter, the contributions of both metrics are reflected. Initial experiments also reveal that relying solely on for data filtering increases hallucinations, as it can lead to mistakenly removing data with high ; whereas using alone does reduce hallucinations, but is not as effective as and will bring greater recall loss. Balancing both metrics yields the most desirable outcomes.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) often suffer from multimodal hallucinations, wherein they may create content that is not present in the visual inputs. In this paper, we explore a new angle of this issue: overly detailed training data hinders the model’s ability to timely terminate generation, leading to continued outputs beyond visual perception limits. By investigating how the model decides to terminate generation with EOS, the special end-of-sentence token, we find that the model assesses the completeness of the entire sequence by comparing the generated text with the image. This observation suggests that the model possesses an inherent potential of making proper EOS decisions based on its visual perception to avoid overly lengthy outputs. To take advantage of such potential, we explore two methods to mitigate multimodal hallucinations: a training objective that enables the model to reduce hallucinations by learning from regular instruction data, and a data filtering strategy to prevent harmful training data from exacerbating model hallucinations. Both methods significantly improve the hallucination performance of LVLMs, without requiring any additional data or knowledge.111https://github.com/yuezih/less-is-more
@@
Zihao Yue Renmin University of China yzihao@ruc.edu.cn Liang Zhang Renmin University of China zhangliang00@ruc.edu.cn Qin Jin††thanks: Corresponding Author. Renmin University of China qjin@ruc.edu.cn
@@ 1 Introduction
Ever since Large Vision-Language Models (LVLMs) Yin et al. (2023a) were achieved through bridging vision encoders with Large Language Models (LLMs) Brown et al. (2020); Zhao et al. (2023a), they have been plagued by the problem of multimodal hallucinations, i.e., their text outputs may include unfaithful content to the visual inputs, such as non-existent objects Rohrbach et al. (2018), which greatly harms the reliability of their applications. Extensive research has shed light on the origins of multimodal hallucinations, including the inability of vision encoders to represent fine-grained visual details Jiang et al. (2023b); Tong et al. (2024), model reliance on inherent parametric knowledge such as language priors and statistical biases Leng et al. (2023); Zhou et al. (2023), and pervasive hallucinations in the training data itself Yu et al. (2023a); Liu et al. (2023a). In response to these insights, a variety of strategies have been proposed to mitigate hallucinations in LVLMs Yin et al. (2023b); Zhai et al. (2023); Zhao et al. (2023b).
Although significant progress has been made, in this paper, we highlight a crucial but often overlooked source of hallucinations: the excessively detailed training data. For example, in the detailed image captioning task, the caption data for an image typically integrates rich visual semantics from multiple human annotations or vision expert models, and is rewritten into lengthy paragraphs by LLMs Liu et al. (2023c), as shown in Fig. 1. These training data, while high-quality and meeting our expectations for detail, may exceed the visual perception capability of LVLMs, especially for subtle image features such as small or easily confusable objects. When trained with such data, in an attempt to fit the detail level and length distribution of ground truth captions, the model may risk expressing details that it cannot discern from the image, and therefore exhibit hallucinations.
@@
To unlock such potential of models, we explore two approaches to enhance the model for better EOS decisions. (1) A learning objective for model training, termed Selective EOS Supervision. Simply modified from the Maximum Likelihood Estimation (MLE), it enables the model to mitigate hallucinations through learning from regular instruction data. It is applicable both for further training to reduce hallucinations in existing models, and for initial instruction tuning Ouyang et al. (2022) to alleviate the onset of hallucinations. Specifically, by briefly further training on the original instruction data, the sentence-level and instance-level hallucinations of LLaVA-1.5 Liu et al. (2023b) are reduced by 26% and 27%, respectively. (2) A data filtering strategy based on Scoring EOS Supervision, to eliminate harmful training data that can impair the model’s ability to end sequences. We design two metrics to assess the positive and negative impact of data on the model’s EOS tendency, and combine them to rank and filter the training data. Experimental results show that removing a small portion of the data can significantly reduce the hallucinations of models trained on it. These findings further validate our hypothesis and provide simple yet effective solutions for mitigating multimodal hallucinations in LVLMs.
@@ 2 EOS Decision
In this section, we delve into how LVLMs reach EOS decisions. Specifically, in Section 2.1, we analyze the contextual information that the model relies on to predict ; in Section 2.2, we explore how the model adjusts its tendency to terminate generation with the multimodal input. Corresponding findings are further discussed in Section 2.3.
@@ 2.1 Information Basis of EOS Decision
where denotes the self-attention score matrix of the language model, means element-wise product, and reflects the significance of the information flow from the -th token to the -th token. We compare the information flow patterns of EOS predictions and non-EOS predictions to elucidate the information basis of EOS decisions.
@@ 3.1 Selective EOS Supervision for Training
The instruction tuning of LVLMs typically utilizes Maximum Likelihood Estimation (MLE) as the training objective. Given the visual content and previous tokens , the model predicts a probability distribution over the vocabulary to determine the next word, where represents the probability of the -th word in . The model parameter is optimized to maximize the likelihood of the label word indexed , with the loss function defined as:
@@
Row | Model | Method | Length | CHAIRS | CHAIRI | Recall | Faith | FaithS |
@@ 3.1.1 Experimental Settings
Models and datasets. Our training objective can be applied to any LVLMs with an EOS token and optimized by MLE. As a representative, we select two widely used open-source LVLMs, LLaVA Liu et al. (2023c) and MiniGPT Zhu et al. (2023) series. Among them, LLaVA, LLaVA-1.5 Liu et al. (2023b), and MiniGPT-v2 Chen et al. (2023a) are trained with data recipes that include the LLaVA-Instruction dataset. Thus, we validate our method with these models by fine-tuning them on LLaVA-Instruction. Additionally, our experiment results show that a smaller subset of LLaVA-Instruction-150K which contains detailed captions, Detail23K, has a similar effect but brings significant computational efficiency. Thus, most of our experiments are conducted with Detail23K. If not specified, models undergo one epoch of training with LoRA Hu et al. (2022). Other training details remain consistent with the models’ official documentation.
@@
Evaluation. Following previous works Huang et al. (2023); Yu et al. (2023a), we evaluate model hallucination with Caption Hallucination Assessment with Image Relevance (CHAIR) Rohrbach et al. (2018), a framework that quantifies object hallucination in image captions by comparing generated objects to the ground truth objects. The sentence-level score, CHAIRS, represents the proportion of captions that contain hallucinations, and the instance-level score, CHAIRI, denotes the frequency of hallucinated objects relative to all mentioned objects by the model. In addition, we measure an object recall to evaluate the semantic richness of generated captions. Our CHAIR tests involve 500 images randomly chosen from the MSCOCO validation set Lin et al. (2014). We also adopt another metric FaithScore Jing et al. (2023) to evaluate caption hallucination. It verifies the consistency between atomic facts in the caption and the input image with LLMs and visual expert models, for which we employ ChatGPT OpenAI (2023) and OFA Wang et al. (2022). It also provides a sentence-level score, FaithScoreS.
@@
Baselines. Since our method facilitates models to timely terminate generation, which often results in shorter responses, we incorporate baselines that simply reduce the generation length, including sequence truncating and decoding with a length penalty. The truncating method keeps only the initial of words in each caption, and the decoding method adopts an exponential length penalty to adjust the score of the EOS token during generation. Varying the truncating proportion or the length penalty factor leads to different generation length, and affects both hallucination and recall performance. Additionally, we include two recently proposed plug-in methods: (1) Visual Contrastive Decoding (VCD) Leng et al. (2023), which contrasts the output distributions derived from the original and noisy visual inputs, to reduce the influence of the model’s parametric knowledge. (2) Over-Trust Penalty and Retrospection-Allocation (OPERA) Huang et al. (2023), a decoding strategy that penalizes the model’s over-reliance on certain tokens and allows roll-back when needed. We test VCD at different noise steps of 200, 500, 700, and 999, and report the optimal results. For OPERA, our implementation follows the suggestions in their released code, including two hyperparameter configurations for standard and fast inference.
@@ 3.1.2 Results
Versus the original model. As shown in Table 1, after a single training epoch on the detailed caption subset, Detail23K, using our learning objective, all models tend to produce shorter captions and notably reduce hallucinations at both the sentence and instance levels ( vs , vs , etc.). This improvement is even more significant when using the full 150K instruction data ( vs ), resulting in a 26.4% and 26.6% decrease in CHAIRS and CHAIRI of LLaVA-1.5 (7b), respectively. While our method leads to some decrease in recall (e.g., 1.8% of LLaVA-1.5 (7b) and 2.6% of LLaVA-1.5 (13b)), we view this as a beneficial compromise since the models become more conservative and less likely to “guess” uncertain visual content. More analysis can be found in Section B.3.
Versus baselines. As demonstrated in Fig. 5, the truncating and length-penalty decoding baselines, with varying length-controlling configurations, effectively reduce hallucinations at the cost of Recall. However, these methods fall short of ours for either being less effective in alleviating hallucinations or resulting in more significant recall loss. Our method also outperforms existing methods, i.e., VCD and OPERA, as shown in Table 1. Our proposed method does not require additional data construction or other expert models. Furthermore, unlike decoding methods, it does not slow down inference and remains technically compatible with various decoding strategies. Therefore, it presents a viable, practical supplement or alternative to current methods.
@@
Loss | Length | CHAIRS | CHAIRI | Recall |
@@ 4 Related Works
Hallucination Origins. Investigations on the causes of multimodal hallucination in LVLMs identify three main factors: (1) Limited visual representations. For example, the visual encoders commonly employed in LVLMs depict abstract features while struggling to capture fine-grained visual details Jiang et al. (2023b); Tong et al. (2024); Wang et al. (2023a). Jiang et al. (2023a) observe a modality gap persists between visual and textual features, despite vision-language alignment. (2) Models’ over-reliance on parametric knowledge, such as statistical biases and language priors, rather than on visual evidence Zhou et al. (2023); Leng et al. (2023); Liu et al. (2023a); Zhai et al. (2023); Guan et al. (2023). (3) Inferior data for instruction tuning. This includes insufficient visual supervision Chen et al. (2023b), a lack of positive/negative human feedback Yu et al. (2023b), and the presence of hallucinations within the training data Yu et al. (2023a); Liu et al. (2023a). Our paper identifies a new source of hallucinations: overly detailed training data hinders the model’s inherent EOS decision ability, further enriching existing explanations.
@@ 5 Conclusion
This work focuses on reducing hallucinations in large multimodal models to enhance their reliability and trustworthiness. We have carefully considered the ethical implications of our work and anticipate no significant ethical concerns. This work was carried out using publicly available and commonly used data and models, and our findings may inherit the biases and limitations carried in these resources.
Acknowledgements
We thank all reviewers for their insightful comments and suggestions. This work was partially supported by the Beijing Natural Science Foundation (No. L233008), the National Natural Science Foundation of China (No. 62072462), and the Outstanding Innovative Talents Cultivation Funded Programs 2023 of Renmin University of China.
@@ B.2 Context Manipulation
Method | #Halluc. | #Correct | Halluc. Rate↑ |
Ours (w/ Inst.) | 263 | 104 | 71.7% |
Ours (w/ Cap.) | 244 | 93 | 72.4% |
Model | #Correct↑ | #Hallucinated↓ |
Original model | 2.45 | 0.90 |
Ours (w/ Cap.) | 2.40 | 0.63 |
Ours (w/ Inst.) | 2.36 | 0.55 |
B.3 Selective EOS Supervision
EOS prediction tendency. In Fig. 9, we illustrate the EOS prediction tendency (average probability) of the LLaVA-1.5 (7b) model during further training on Detail23K. With Selective EOS Supervision proposed in Section 3.1, the model’s tendency to predict EOS rises and stabilizes, while the model optimized by MLE shows no change in this behavior. This suggests that the proposed training objective effectively helps the model regain its capability to timely conclude sequences.
Dissecting omitted content. As our method reduces the generated content to alleviate hallucinations, it is interesting to investigate what is “omitted” by our method from the originally generated captions, specifically, how many “omitted” objects are correct and how many are hallucinations. We extract the generated objects from the outputs of both the original model and our further trained models, using the same technique as in the CHAIR evaluation. Then, we focus on these objects that are mentioned by the original model but not by our models, which are “omitted” from the original captions. As the results of the Halluc. Rate (hallucinated object rate of omission) in Table 5 shows, nearly 3/4 of the “omitted” objects are hallucinations, implying that such an omission is beneficial.
Furthermore, we analyze the average counts of correct and hallucinated objects in the model generation, as a supplement to the CHAIR metrics, to demonstrate more comprehensively how our method impacts the quality of model generation. As shown in Table 6, our method reduces hallucinations while largely preserving the correct content.
@@ B.3 Selective EOS Supervision
LLaVA-1.5 (7b) | Baseline | 1,516.1 | 348.2 | 85.9 | 86.9 | 94.0 | 79.1 |
MLE | 1,470.9 | 372.5 | 86.1 | 87.0 | 93.6 | 79.7 | |
Ours | 1,490.4 | 367.9 | 86.0 | 86.8 | 93.5 | 79.5 | |
@@
B.4 Scoring EOS Supervision
Data Score Distributions. In Section 3.2, we discuss two metrics, and , which are summed over positions labeled as EOS and non-EOS, respectively. A natural concern is that the number of non-EOS positions far exceeds that of EOS positions, raising the question of whether the combined might be dominated by . To clarify this, we examine the score distributions within the LLaVA-Instruction-150K dataset. As illustrated in Fig. 10, the value distributions of and are comparable in magnitude, and the distribution is approximately normal with zero mean. This indicates a balance between and with neither metric dominating, and the top scores necessitate both high and low . Thus, by maintaining a straightforward formulation of without introducing a balancing hyperparameter, the contributions of both metrics are reflected. Initial experiments also reveal that relying solely on for data filtering increases hallucinations, as it can lead to mistakenly removing data with high ; whereas using alone does reduce hallucinations, but is not as effective as and will bring greater recall loss. Balancing both metrics yields the most desirable outcomes.